OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5

OpenStudy (anonymous):

what do you want solved?

OpenStudy (across):

Expand both the LHS and the RHS.

OpenStudy (anonymous):

Is it 6?

OpenStudy (anonymous):

When i worked through the question thats what i got. @across

OpenStudy (across):

Let's see:\[-3(x+3)=-3(x+1)-5\\-3(6+3)=-3(6+1)-5\\-3(9)=-3(7)-5\\-27=-21-5\\-27\neq-26\]I don't think so. :-/

OpenStudy (anonymous):

Why do you put 6 as x?

OpenStudy (anonymous):

IM trying to solve for x

OpenStudy (across):

I'll solve it for you. You're given\[-3(x+3)=-3(x+1)-5.\]Simplifying yields\[-3x-9=-3x-3-5\\-9=-8.\]Because the result is absurd, this means that the lines are parallel and hence never intersect.

OpenStudy (anonymous):

So what does x equal ?

OpenStudy (anonymous):

Is there no solution for it ?

OpenStudy (across):

There is no solution.

OpenStudy (anonymous):

Okay i understand